What volume does 6.76 × 1021 molecules of N2 occupy at STP?

1
Expert's answer
2021-03-29T06:09:34-0400

1 mole of N2 = 6.022 × 10^ 23

= 6.76 × 10 ^21/ 6.022 × 10^ 23

= 0.01moles of N2

1 mole of any gas ocuppy 22.4 L

= 22.4 × 0.1

= 2.24L